OpenStudy (kanwal32):

In quadratic equation ax^2+bx+c=0 ... (sequence and series)

OpenStudy (kanwal32):

IN quadratic equation ax^2+bx+c=0. If alpha,beta are the roots of the equation d=b^2-4ac and alpha+beta,alpha^2+beta^2,alpha^3+beta^3 are in GP them

OpenStudy (kanwal32):

A)d is not equal to zero B)bd=0 C=cd=0 D)d=0

OpenStudy (kanwal32):

@dan815 @ganeshie8 @ikram002p @Miracrown

OpenStudy (kanwal32):

@nincompoop

OpenStudy (kanwal32):

IIT advanced question

OpenStudy (kanwal32):

@Kainui

ganeshie8 (ganeshie8):

establish below first : \(\large \alpha + \beta = -\dfrac{b}{a}\) \(\large \alpha^2 + \beta^2 = \dfrac{b^2}{a^2} - 2\dfrac{c}{a}\) \(\large \alpha^3 + \beta^3 = -\dfrac{b}{a}\left( \dfrac{b^2}{a^2} - 3\dfrac{c}{a}\right)\)

OpenStudy (kanwal32):

yes

OpenStudy (kanwal32):

i established these

ganeshie8 (ganeshie8):

after that use the geometric mean to get a relation in a,b,c simplifying it yields \(b^2-4ac = 0\) or, \(d = 0\)

OpenStudy (kanwal32):

i got the same answer but the answer is given as cd=0

OpenStudy (kanwal32):

\[does \Delta\] or d makes any difference to u

OpenStudy (kanwal32):

@ganeshie8

ganeshie8 (ganeshie8):

can you take a screenshot and attach if possible - just for clarity

OpenStudy (kanwal32):

OK